Grade 11Mechanics

The position of a particle moving rectilinearly is given by t =t^3-3t^2-10. Find the distance travelled by the particle in the first four seconds starting from t=0

saheed ahmed
10 Years agoGrade 11

1 Answer

Mudit
9 Years ago
x=t3-3t2-10
dist. travelled in4 sec=43-3(4)2-10=64-48-10=6unit
Plz approve my answer if ufind it helpful.:)